Find the unknown quantities a. P1(-6, y), P2 (8, 4) and P1P2 = 13
Answer #1

P. (-6,y) P2 (8,4) P, P₂ = 13 distance formula = 562-2)²+ (yz-y, J² 13 = J [8-(-6)]” + [4-y] squaring both sides 169= (14)² +